How fast in rpm must a centrifuge rotate if a particle 7.0cm from the axis of rotation is to experience an acceleration of 200000 g's?
Answer #1

centrifuge acceleration = r\omega2
\omega in rad/s
here acceleration =200000g m/s2   (g=9.8 m/s2 )

2 * 105 * 9.8 = 0.7 * 10-2 *\omega2
\omega2 = 280000000​

\omega= 16733.2 rad/s

\omega=159790.29 rpm